OpenStudy (anonymous):

5x-3-(2-3x) please help and show work?? Thank you!

OpenStudy (anonymous):

hey... isnt this basic algebra

OpenStudy (anonymous):

...maybe xD but idk the first step!!

OpenStudy (anonymous):

expand the brackets. first

OpenStudy (anonymous):

think of it as 5x-3-1*(2-3x)

OpenStudy (anonymous):

so 5x-2(2-3x) then distribute 2 times 2 and 2 times -3x ?

OpenStudy (anonymous):

so. 5x-3-(2-3x) =5x-3-1(2-3x) =5x-3+(-1*2+-1*-3x) =5x-3+(-2+3x) =5x-3-2+3x =8x-5

OpenStudy (anonymous):

\[5x-3-(2-3x)\]distribute the minus sign \[=5x-3-2+3x\] like terms are \(5x\) and \(3x\) which add up to \(8x\) also \(-3-2=-5\) you get \(8x-5\)